# Asymmetric Proximal Policy Optimization: mini-critics boost LLM reasoning

### v1

@@ C.2 Hyperparameters
C.3 Prompt
[…]
@@ C.3 Prompt
Appendix D The relationship between value std and state information quantity
Specifically, for the training scenarios of 8b actors and two 0.6b critics, we use the value-std corresponding to the global state and the median of the gradient magnitude to categorize the states into four types. Namely, large gradient & large value std, large gradient & small value std, small gradient & large value std, small gradient & small value std. The results in Figure 10 (Left) show that the vast majority of states are classified into the categories of large gradient & large value std and small gradient & small value std, thereby empirically proving the positive relationship between value std and the learning value (information quantity) of the state.
Appendix E Visualization of word clouds

### v2

@@ A.3 Hyperparameters
Appendix B The relationship between value std and state information quantity
Specifically, for the training scenarios of 8b actors and two 0.6b critics, we use the value-std corresponding to the global state and the median of the gradient magnitude to categorize the states into four types. Namely, large gradient & large value std, large gradient & small value std, small gradient & large value std, small gradient & small value std.
@@ Appendix B The relationship between value std and state information quantity
The results in Figure 11 (Left) show that the vast majority of states are classified into the categories of large gradient & large value std and small gradient & small value std, thereby empirically proving the positive relationship between value std and the learning value (information quantity) of the state.
Appendix C Visualization of word clouds